# CPRM: A LLM-based Continual Pre-training Framework for Relevance Modeling in Commercial Search

### v2

@@ 1 Introduction
To address the above problems, we investigate a Continual Pre-training approach of LLMs for Relevance Modeling, CPRM for short. Initially, we introduce a pre-training technique using pairs of queries and multi-field item as inputs. This method enables the LLMs to explicitly model the semantic representations between queries and items, thus bridging the semantic gap between them. Subsequently, we collect sets of semantically similar queries and items based on user click logs, then further refine these sets through semantic modeling to filter out semantically irrelevant cases. Following that, we reorder these refined sets according to semantic similarity and ultimately construct in-context pre-training instances via prompting techniques. Utilizing this approach to data construction, LLMs are able to make better predictions within such contexts during the training process, which benefits the efficient learning current domain knowledge for LLMs. Lastly, we employ ChatGPT to conduct reading comprehension on structured item data, facilitating the generation of relevant domain knowledge for pre-training, which can be considered as a secondary development and exploitation of the item. More specifically, we leverage ChatGPT to summarize and paraphrase item to produce fluent domain knowledge, while also guiding ChatGPT to produce background information related to the items. Additionally, we prompt ChatGPT to create diverse queries and provide further reasons for their generation. In summary, the contributions of this paper are as follows:
@@ 2 Related Work
In recent years, pre-trained models like BERT (Devlin et al., 2019) has show its superiority on natural language understanding (NLU) tasks. Consequently, both representation-based and interaction-based methods have begun leveraging the capabilities of these pre-trained models for semantic modeling. Most recently, LLMs like GPT-3 (Brown et al., 2020), GLM (Du et al., 2022), LLaMA (Touvron et al., 2023a), Qwen (Bai et al., 2023; Yang et al., 2024) pre-trained extensive volumes of data with numerous parameters have garnered significant performance in language understanding, generation and reasoning tasks. Compared to traditional pre-trained models like BERT (Devlin et al., 2019), LLMs possess significant advantages in both the scale of pre-training data and the quantity of model parameters, leading to their evident superiority in performance across a variety of downstream tasks. Recent research work (Sun et al., 2023; Spatharioti et al., 2023; Zhu et al., 2024) indicates that combining LLMs with downstream applications presents significant potential, LLMs can achieve competitive or even superior results compared to traditional supervised methods on information retrieval benchmarks. Some research leverage LLMs for relevance modeling in search engines, adopting approaches such as behavior-augmented (Chen et al., 2023, 2024) or robust learning (Liu et al., 2024) to improve the capability of relevance modeling. Our work mainly focuses on enhancing LLMs from the perspective of continual pre-training for relevance modeling.
@@
LLMs are pre-trained on a wide variety of data sources (Brown et al., 2020; Du et al., 2022; Touvron et al., 2023a, b) without pay more attention on specific domains, resulting in a lack of domain knowledge. On the other hand, the pre-training phase of LLMs is task-agnostic (Brown et al., 2020), making it difficult to direct connect with downstream tasks. This also means we can’t easily customize the pre-training process to better fit those tasks (Min et al., 2022; Gu et al., 2023). Previous work of injecting domain knowledge involves continued training of pre-trained models on domain-specific data (Gururangan et al., 2020; Shi et al., 2023), as well as incorporating knowledge graphs (Liu et al., 2019a) or selectively masking important information (Gu et al., 2020; Xu et al., 2023; Sanyal et al., 2023; Zhou et al., 2023). Another line of research aims to enhance the pre-training for downstream tasks, which simple concatenate relevant documents together for in-context pre-training (Min et al., 2022; Gu et al., 2023; Shi et al., 2024). However, these approaches assume that downstream tasks contain only a single domain representation, neglecting the possibility of there being multiple or more. For instance, in commercial search relevance tasks, queries and items belong to two distinct domains with substantial differences. Our research work involves injecting domain knowledge and conducting in-context pre-training simultaneously, while being able to establish a connection between the two domains.
@@ 3.1 Domain Knowledge Enhancement (DKE)
$\mathcal{L}(\theta)=\mathop{\textrm{min}}\limits_{\theta}\alpha\mathcal{L}_{\textrm{t-MLM}}(\theta)+(1-\alpha)\mathcal{L}_{\textrm{s-MLM}}(\theta).$ (1)
@@ 3.2 In-Context Pre-training (ICP)
$Sim(Query,I_{k})<\sigma,\quad k\in[1,N]$ (2)
@@ 3.3 Reading Comprehension Distillation (RCD)
We employ ChatGPT for reading comprehension on items, with the prompt design shown in Figure 3. Assume that in a mini-app search scenario, we need to provide the mini-app’s structured information like title, keywords, category, and description, and utilizing prompt template instructions to invoke ChatGPT. This generates the reading comprehension pre-training instances. Why design the prompt in this way? We have several reasons for this design choice. Firstly, item text is structured data and difficult to utilize, lacking in relevant background knowledge. Through summarizing and rephrasing with Prompt 1, fluent domain knowledge can be generated. Additionally, understanding and analyzing items can instruct ChatGPT in generating relevant background knowledge. Secondly, by using Prompt 2 and Prompt 3 guide ChatGPT to generate related and diversified queries, enriching the supply of suitable queries. We also instruct ChatGPT to provide further explanations for the generated queries. This approach not only facilitates the generation of relevant domain knowledge but also allows downstream models to significantly improve their understanding and handling of the item when utilizing these data. Employing ChatGPT for reading comprehension on items can be considered a secondary development and utilization of item data, enriching the domain knowledge further. Pre-training LLMs on the above data can also be seen as a process of knowledge transfer from ChatGPT to LLMs.
@@ 4.1 Experimental Settings
Pre-training Dataset. We utilize the real-world mini-app search scenario data for verification. The pre-training data includes three parts: DKE data (4M), ICP data (4M) and RCD data (500K). The first part is sampled from the mini-app search scenarios and consist of structured items containing multiple fields. For top 500K most frequently visited items, we sample 5 top queries based on click logs for each item, which are then concatenated with multi-field item to serve as pre-training examples for adapting relevance tasks. The second part is in-context pre-training data, where we construct these examples based on the real-world search click logs using the method described in Section 3.2, and subsequently randomly sample 4M from them. The third part is reading comprehension data, for which we utilize ChatGPT to perform reading comprehension on item data.
Relevance Dataset. The SFT data consists of three parts: train set (625K), valid set (35K) and test set (35K). These data are sourced from real mini-app search results and then are generated through manual annotation. The human-annotated data for relevance tasks are in format of triples <Query, Item, Label>, the data statistics as shown in Table 1. The annotated data have only two levels of relevance, #Related and #Not_Related denote label 1 and 0 respectively.
Evaluation Metrics. We employ three widely used metrics Acc., F1 and AUC to evaluate model performance, with higher values indicating better performance. Note that AUC serves as the most important metric in relevance tasks while the others provide auxiliary supports for our analysis.
@@
Dataset #Sample #Query #Item #Related #Not_Related
@@
11 ChatGPT (zero-shot) 62.3 - 69.3 - 60.8 -
12 GPT-4 (zero-shot) 65.0 - 71.3 - 63.2 -
@@ 4.2 Baselines
  • ChatGPT & GPT-4 (OpenAI, 2024) are the SOTA closed-source LLMs. They cannot perform SFT and employ the direct generation approach for relevance task evaluation.
@@ 4.3 Offline Experimental Results
Performance Comparison. Table 2 presents the performance of different baselines and various continual pre-training models on the relevance task. From the experimental results, GLM demonstrates strong competitiveness, achieving superior performance even with similar parameter numbers compared to BERT-Large (line 7 vs. line 4). We also conducts experiments on the GLM of various parameter sizes, and the results show that as the model size increases, its performance gradually improves. However, with further increases in model size, the performance gains become progressively smaller. Specifically, GLM-10B achieves only a 0.21% improvement in AUC over GLM-5B (line 10 vs. line 9). Compared to other latest LLMs such as Qwen2, our GLM also shows impressive performance at a similar parameter scale (line 7 vs. line 5, line 8 vs. line 4). ChatGPT and GPT-4 performed poorly compared to other SFT-based baseline systems; this is because these models are more suited for NLG tasks and are not capable of supporting SFT. Additionally, the task data is biased towards a specific domain. We conducts continued pre-training experiments on GLM-2B, and the experimental results demonstrate that all three different methods result in performance improvements compared to the baseline model. Notably, the DKE and ICP methods achieve significant performance enhancements, with respective gains of 1.11% and 1.54% in AUC. This is because both methods are constructed based on domain-specific data and jointly training semantically related queries or items can further enhance model performance. The experimental results also indicate that integrating different continued pre-training methods can further strength model performance (line 17 and line 18), with the combination of all three methods leading to the greatest performance gain, making it comparable to that of GLM-10B (line 18 vs. line 10). Our CPRM model achieves the highest F1 score (94.42%) among all models compared.

### v5

@@ 1 Introduction
To address the above problems, we investigate a Continual Pre-training approach of LLMs for Relevance Modeling, CPRM for short. Initially, we introduce a pre-training technique using pairs of queries and multi-field item as inputs. This method enables the LLMs to explicitly model the semantic representations between queries and items, thus bridging the semantic gap between them. Subsequently, we collect sets of semantically similar queries and items based on user click logs, then further refine these sets through semantic modeling to filter out semantically irrelevant cases. Following that, we reorder these refined sets according to semantic similarity and ultimately construct in-context pre-training instances via prompting techniques. Utilizing this approach to data construction, LLMs are able to make better predictions within such contexts during the training process, which benefits the efficient learning current domain knowledge for LLMs. Lastly, we employ an LLM with larger parameters (called teacher LLM) to conduct reading comprehension on structured item data, facilitating the generation of relevant domain knowledge for pre-training, which can be considered as a secondary development and exploitation of the item. More specifically, we leverage teacher LLM to summarize and paraphrase item to produce fluent domain knowledge, while also guiding teacher LLM to produce background information related to the items. Additionally, we prompt teacher LLM to create diverse queries and provide further reasons for their generation. In summary, the contributions of this paper are as follows:
@@ 2 Related Work
In recent years, pre-trained models like BERT (Devlin et al., 2019) has show its superiority on natural language understanding (NLU) tasks. Consequently, both representation-based and interaction-based methods have begun leveraging the capabilities of these pre-trained models for semantic modeling. Most recently, LLMs like GPT-3 (Brown et al., 2020), GLM (Du et al., 2022), LLaMA (Touvron et al., 2023a), Qwen (Bai et al., 2023; Yang et al., 2024) pre-trained extensive volumes of data with numerous parameters have garnered significant performance in language understanding, generation and reasoning tasks. Compared to traditional pre-trained models like BERT (Devlin et al., 2019), LLMs possess significant advantages in both the scale of pre-training data and the quantity of model parameters, leading to their evident superiority in performance across a variety of downstream tasks. Recent research work (Sun et al., 2023; Spatharioti et al., 2023; Zhu et al., 2024) indicates that combining LLMs with downstream applications presents significant potential, LLMs can achieve competitive or even superior results compared to traditional supervised methods on information retrieval benchmarks. Some research leverage LLMs for relevance modeling in search engines, adopting approaches such as behavior-augmented (Chen et al., 2023, 2024) or robust learning (Liu et al., 2024) to improve the capability of relevance modeling. Our work mainly focuses on enhancing LLMs from the perspective of continual pre-training for relevance modeling. LLMs are pre-trained on a wide variety of data sources (Brown et al., 2020; Du et al., 2022; Touvron et al., 2023a, b) without pay more attention on specific domains, resulting in a lack of domain knowledge. On the other hand, the pre-training phase of LLMs is task-agnostic (Brown et al., 2020), making it difficult to direct connect with downstream tasks. This also means we can’t easily customize the pre-training process to better fit those tasks (Min et al., 2022; Gu et al., 2023). Previous work of injecting domain knowledge involves continued training of pre-trained models on domain-specific data (Gururangan et al., 2020; Shi et al., 2023), as well as incorporating knowledge graphs (Liu et al., 2019a) or selectively masking important information (Gu et al., 2020; Xu et al., 2023; Sanyal et al., 2023; Zhou et al., 2023). Another line of research aims to enhance the pre-training for downstream tasks, which simple concatenate relevant documents together for in-context pre-training (Min et al., 2022; Gu et al., 2023; Shi et al., 2024). However, these approaches assume that downstream tasks contain only a single domain representation, neglecting the possibility of there being multiple or more. For instance, in commercial search relevance tasks, queries and items belong to two distinct domains with substantial differences. Our research work involves injecting domain knowledge and conducting in-context pre-training simultaneously, while being able to establish a connection between the two domains.
3 Problem Formulation
[…]
$\textrm{P}(Q,I)=Is\ [Q]\ and\ [I]\ related?\ \ [\textrm{Mask}]$ (1)
[…]
where relevance label $y\in\{0,1\}$ can be associated with a verbalizer (e.g., “no” or “yes”) from the vocabulary of LLMs to represent “irrelevant” and “relevant” between $Q$ and $I$ respectively. To enable the adaptation of general LLMs to the relevance modeling task, SFT operation is selected for training with the cross-entropy loss function. Consequently, the relevance degree could be given from LLMs for subsequent applications in search scenarios.
@@ 4.1 Domain Knowledge Enhancement (DKE)
$\mathcal{L}(\theta)=\mathop{\textrm{min}}\limits_{\theta}\alpha\mathcal{L}_{\textrm{t-MLM}}(\theta)+(1-\alpha)\mathcal{L}_{\textrm{s-MLM}}(\theta),$ (3)
@@ 4.2 In-Context Pre-training (ICP)
$Sim(Query,I_{k})<\sigma,\quad\textrm{for}\ k\in[1,N],$ (4)
@@ 4.3 Reading Comprehension Distillation (RCD)
We employ the teacher LLM for reading comprehension on items, with the prompt design shown in Figure 3. Assume that in a mini-app search scenario, we need to provide the mini-app’s structured information like title, keywords, category, and description, and utilizing prompt template instructions to invoke teacher LLM. This generates the reading comprehension pre-training instances.
Why design the prompt in this way? We have several reasons for this design choice. Firstly, item text is structured data and difficult to utilize, lacking in relevant background knowledge. Through summarizing and rephrasing with Prompt 1, fluent domain knowledge can be generated. Additionally, understanding and analyzing items can instruct teacher LLM in generating relevant background knowledge. Secondly, by using Prompt 2 and Prompt 3 guide teacher LLM to generate related and diversified queries, enriching the supply of suitable queries. We also instruct teacher LLM to provide further explanations for the generated queries. This approach not only facilitates the generation of relevant domain knowledge but also allows downstream models to significantly improve their understanding and handling of the item when utilizing these data. Employing teacher LLM for reading comprehension on items can be considered a secondary development and utilization of item data, enriching the domain knowledge further. Pre-training LLMs on the above data can also be seen as a process of knowledge transfer from teacher LLM to LLMs.
@@ 5.1 Experimental Settings
Dataset & Evaluation Metrics. We utilize the real-world mini-app search scenario data for verification. The pre-training data includes three parts: DKE data (4M), ICP data (4M) and RCD data (500K). The first part is sampled from the mini-app search scenarios and consist of structured items containing multiple fields. For top 500K most frequently visited items, we sample 5 top queries based on click logs for each item, which are then concatenated with multi-field item to serve as pre-training examples for adapting relevance tasks. The second part is in-context pre-training data, where we construct these examples based on the real-world search click logs using the method described in Section 4.2, and subsequently randomly sample 4M from them. The third part is reading comprehension data, for which we utilize teacher LLM (e.g. Qwen2-72B (Yang et al., 2024)) to perform reading comprehension on item data. The SFT data consists of three parts: train set (625K), valid set (35K) and test set (35K). These data are sourced from real mini-app search results and then are generated through manual annotation. The human-annotated data for relevance tasks are in format of triples <Query, Item, Label>, the data statistics as shown in Table 1. The annotated data have only two levels of relevance: “#Relevant” and “#Irrelevant”. For evaluation, we employ three widely used metrics Acc., F1 and AUC to evaluate model performance, with higher values indicating better performance. Note that AUC serves as the most important metric in relevance tasks while the others provide auxiliary supports for our analysis.
@@
Dataset #Sample #Query #Item #Relevant #Irrelevant
@@
11 ChatGPT (+8-shot) 62.93 - 59.78 - 64.88 -
12 GPT-4 (+8-shot) 61.89 - 67.44 - 60.83 -
@@ 5.2 Baselines
  • ChatGPT [^2] The version is GPT-3.5-Turbo. & GPT-4 (OpenAI, 2024) are the SOTA closed-source LLMs. We employ the direct generation approach for relevance task evaluation.
@@ 5.3 Offline Experimental Results
Performance Comparison. Table 2 presents the performance of different baselines and various continual pre-training models on the relevance task. From the experimental results, GLM demonstrates strong competitiveness, achieving superior performance even with similar parameter numbers compared to BERT-Large (line 7 vs. line 4). We also conducts experiments on the GLM of various parameter sizes, and the results show that as the model size increases, its performance gradually improves. However, with further increases in model size, the performance gains become progressively smaller. Specifically, GLM-10B achieves only a 0.21% improvement in AUC over GLM-5B (line 10 vs. line 9). Compared to other latest LLMs such as Qwen2, our GLM also shows impressive performance at a similar parameter scale (line 7 vs. line 5, line 8 vs. line 6). ChatGPT & GPT-4 performed poorly compared to other SFT-based baseline systems; this is because the task data belongs to a proprietary domain, and models without SFT operations have relatively poor discriminative ability. We conducts continued pre-training experiments on GLM-2B, and the experimental results demonstrate that all three different methods result in performance improvements compared to the baseline model. Notably, the DKE and ICP methods achieve significant performance enhancements, with respective gains of 1.11% and 1.54% in AUC. This is because both methods are constructed based on domain-specific data and jointly training semantically related queries or items can further enhance model performance. The experimental results also indicate that integrating different continued pre-training methods can further strength model performance (line 17 and line 18), with the combination of all three methods leading to the greatest performance gain, making it comparable to that of GLM-10B (line 18 vs. line 10). Our CPRM model achieves the highest F1 score (94.42%) among all models compared.
@@ 8 Conclusion
Appendix A Case Study
As shown in Figure  7, we provide several cases to compare the relevance output results between the baseline (GLM-2B) and CPRM models. From these cases, we can observe that the CPRM method is able to supplement additional domain knowledge to correct erroneous prediction results. Furthermore, CPRM demonstrates a stronger understanding of long and complex queries (query length greater than 15) compared to the baseline. The SFT data format can be referred to in Figure  8.
Figure 7: Case study.
Figure 8: SFT data examples.